The enthalpy of formation of NH3(g) as per the following reaction is -46.11 kJ/mol

at 298 K.

N2(g) + 3H2 (g) --> 2NH3(g)

Calculate the value of enthalpy of formation of NH3(g) at 100 °C. The Cp

values are given as: N2(g) = 29.12; H2(g) = 28.82; NH3(g) = 35.06. Assume the

given heat capacity values to be temperature independent in the range.


1
Expert's answer
2021-01-15T06:44:58-0500

"\Delta"rCp = 2*35.06 - 3*28.82 - 29.12 = -45.46

"\Delta"fH373К = 1/2(2"\Delta"rH298К + "\Delta"rCp*(373-298)) = 1/2(- 2*46110 - 45.46*75) = - 47814.75 J/mol = - 47.81 kJ/mol